Changes between Initial Version and Version 1 of UC20170612


Ignore:
Timestamp:
Jul 30, 2021, 7:48:52 AM (10 months ago)
Author:
trac
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • UC20170612

    v1 v1  
     1== APO 3.5-m Users Committee Phonecon, 6/12/17 ==
     2
     3Attending:  Nancy Chanover, Ben Williams, Mark Klaene, Scott Anderson, Chip Kobulnicky, John Wisniewski, Rachel Kuzio de Naray, Russet !McMillan, Mary Beth Kaiser, John Bally, Anne Verbiscer, Bill Ketzeback, John Wilson, Jon Holtzman, Leslie Hebb, Joanne Hughes, and Bruce Gillespie
     4
     5----
     6
     7=== User feedback and comments from institutional representatives: ===
     8
     9* Washington - Scott had nothing new to report.
     10* Georgia State - Rachel had nothing new to report.
     11* Seattle - Joanne had nothing new to report.
     12* NMSU - Jon had nothing new to report.
     13* JHU - Mary Beth had nothing new to report.
     14* Oklahoma - John W. had nothing new to report.
     15* Colorado - John B. had nothing new to report.
     16* Wyoming - Chip had nothing new to report.
     17* Virginia - Anne had nothing new to report.
     18* NAPG - Leslie reported having a recent issue involving observing time scheduling, and asked if there could be a general posting of the scheduling policies, including rules for trades and other allocation conventions.  Nancy said she would construct such a message and post it to general distribution.
     19
     20----
     21
     22=== Telescope and Instruments Report: ===
     23
     24Mark noted that contamination on the DIS and echelle detectors is being monitored closely, hoping that we can put off any purging until the shutdown.  Most everything else has been working fairly well of late; the !TripleSpec repair work is discussed separately.  On the DIS and echelle detector contamination, Nancy added that the instruments are currently producing good data — if users feel otherwise they should let her know.  Mary Beth mentioned that it would have been useful to have known in advance that the echelle detector temperature had been raised.  She also feels that the instrument logs are not being adequately kept up to date.  This evoked an extended discussion — Nancy will work with Bill Ketzeback on ensuring that instrument status is correctly reflected in our logs and other postings.  As to the telescope and instrument report, Mark’s more detailed material follows:
     25
     26
     27''3.5-m Telescope and Instruments Highlights, 5/5/17 through 6/6/17''
     28
     290) Overview
     30
     31It has been a wetter than usual May for southern NM, which has not only kept threat of forest fire down as well as actually allowed for some prescribed fire burning.  None of these have affected the observatory.  We did have a couple wind and dust events but nothing compared to a more typical May.  We had a fair amount of site activity hosting; an SDSS workshop, veterans tour, Furman Univ. tour, Discovery TV and JHU training run, and the NAIC visiting instrument.  Also the CSUN Portable Adaptive Optics instrument team is currently on site for a 3rd visit.
     32
     331) Telescope - Telescope was operational with only minimal issues.
     34
     352) Instruments
     36
     37DIS blue and red cameras continue to be operational along with the slitviewer.  No real measureable change to contamination level.
     38
     39!TripleSpec servicing by the UVa team was completed and we expect it to be quite successful once it come back on line.  See agenda item for more information.  We do plan to move it to the observing level soon and after a final vacuum check start cooling next week.
     40
     41Agile has been operational this period.
     42
     43Echelle had been operational with no problems.  The Inter-order Light in red has stayed consistent and good.   In the blue it is still well above our concern point, the downward trend has slowed but still moving slightly downward (worse).
     44
     45NIC-FPS is operational.
     46
     47ARCTIC is operational with no change
     48
     49
     503) ARCSAT - several small improvements and some engineering maintenance were completed this period.  Operation has been fairly routine with occasional typical commercial software issues and limitations.
     51
     52
     53----
     54
     55=== !TripleSpec status update ===
     56
     57John Wilson and his team from UVa were recently at APO to look into the long-standing electronic problems with !TripleSpec.  With the instrument warm and in the lab, they were able to find spurious signals between the controller and the detector, and these were duly fixed.  John doesn’t know for sure that this was the cause of the !TripleSpec readout problems, but he is fairly confident that it was.  Additionally, the UVa team found that the slit wheel motor controller had somehow flipped from 1/2-step to 1/4-step mode, which was also reset to the correct value.  In addition to other minor electronic upgrades, all of the vacuum KF and main O-rings were found to have deteriorated over the past eight years and were replaced, as were the vacuum gauge and positive pressure pop-off.  When the instrument is next under vacuum and cold, the spectrograph focus will need to be checked and may need to be adjusted.  Bill said that !TripleSpec has been on the vacuum pump for a couple of weeks and is now back at the telescope getting re-cabled.  The first tests will be to confirm proper communication with the electronics, then the instrument will be cooled for array testing.  On-sky engineering is planned for !TripleSpec on 28 June.
     58
     59----
     60
     61=== ARCES status update ===
     62
     63Bill said that this topic was adequately covered in the telescope and instrument report above.
     64
     65----
     66
     67=== Q2 scheduling (3.5-m, ARCSAT ): ===
     68
     69Russet said that all the Q2 time has been assigned for the 3.5-m telescope and ARCSAT.
     70
     71----
     72
     73=== Q3 3.5-m scheduling: ===
     74
     75Russet and Ben announced that the Q3 schedule was just issued, with OPEN and DD03 half-nights still available.  The ARCSAT call for proposals will go out today. If the DIS and/or echelle detector contamination gets significantly worse, the August schedule may need to be recast.
     76
     77----
     78
     79=== Open position for Research Scientist / Engineer 1: ===
     80
     81This was an FYI item regarding the posting of a position announcement to replace Jason Lozo, who recently left the project.
     82
     83----
     84
     85=== Debrief from meeting on “Building the Infrastructure for Time-Domain Alert Science in the LSST Era”: ===
     86
     87Nancy reported on her recent attendance at the referenced workshop in Tucson, which depicted the operations concept for the “LSST ecosystem”.  Nancy came away with a better understanding of how APO could potentially participate in support of LSST — she now wants to ascertain how much interest there is within ARC institutions in time-domain astrophysics.  To this end, Ben will head up a group to look into this in more detail, and if others want to participate in this process, they should contact Nancy and Ben.  There is the potential of supporting ZTF next year, and we should try to fully understand how strong our capabilities are for contributing to time-domain science, including the need to interact with the TD special-interests groups.  This led to the following:
     88
     89ACTION ITEM: Users Committee members to query user interest in time-domain science using the 3.5-m telescope, and get back to Nancy and Ben as soon as possible.
     90
     91----
     92
     93=== How best to showcase scientific productivity of 3.5-m telescope at ARC BoG Meeting: ===
     94
     95Nancy brought up the topic of how we highlight our scientific achievements at the annual ARC Board of Governors meeting.  At these meetings, there has been a long-standing tradition of showing a handful of slides and discussing one or two science highlights from each ARC institution.  Nancy wonders if this is the best method to present the scientific productivity of the telescope.  Scott said he thought that these highlights slides were most useful for the board members that are not astronomers (every ARC institution has both an astronomer and an administrator as board members).  John W. thought that the more relevant information would come from comparing the output of the 3.5-m to other comparable telescopes, e.g., counts of refereed publications (this has also been done in the past).  Jon said that we could use our science results summaries as advertising for new students at the ARC universities.  Mary Beth cautioned that we need to be careful because often the annual science highlights are results from not-yet-published research projects.  Mark added that we should try to maintain a balance of reports between the ARC institutions.  Anne suggested also listing the newly trained students in the reports.  Nancy concluded by thanking everyone for their feedback on this; she will continue to mull over this topic.
     96
     97----
     98
     99=== Bruce Gillespie's last meeting: ===
     100
     101With his retirement approaching on July 1, Nancy noted that this was Bruce's final Users Committee telecon.  The members of the Committee all thanked Bruce for his many years of service.  Nancy also expressed an appreciation for his thorough recording of minutes from the UC calls.  Thanks Bruce!  You will be missed.
     102
     103----
     104
     105
     106=== ACTION ITEMS:===
     107
     108'''open from previous meetings:'''
     109None
     110
     111
     112'''new from this meeting:'''
     113* Users Committee members to query user interest in time-domain science using the 3.5-m telescope, and get back to Nancy and Ben as soon as possible.
     114* Nancy will work with Ben and Russet on a web page that lists the telescope scheduling policies to avoid any future misunderstandings.
     115* Nancy will work with Bill to make sure that the Instrument News page reflects the current status information that users are looking for.
     116
     117----
     118
     119=== Next meeting:===
     120The next Users Committee phonecon will be on Monday 10 July at 1030 MDT — the agenda and other materials will be sent to committee members during the preceding week.